# Rademacher’s Theorem for Calderon-Zygmund-type Spaces

### v2

@@ 2 Calderón-Zygmund spaces with Function parameter
The condition $\underline{b}(\phi)>-\frac{d}{p}$ is here to ensure that the spaces are not degenerate. One can easily check that the polynomial appearing in (3) is unique. We set
@@
Then it is straightforward that $(T^{p}_{\phi}(x_{0}),\|\cdot\|_{T^{p}_{\phi}(x_{0})})$ is a Banach space. Let $t^{p}_{\phi,n}(x_{0})$, constitued of functions $f\in L^{p}(\mathbb{R}^{d})$ such that there exists a polynomial $P$ of degree at most $n$ such that
@@
If $\phi(t)=t^{u}$, with $u\in\mathbb{N}_{0}$, then $t^{p}_{\phi,u}(x_{0})$ is the usual space $t^{p}_{u}(x_{0})$ of [17]. Remark that, in this case, the space $T^{p}_{\phi}(x_{0})$ involves a polynomial of degree at most u1u-1 whereas the space $t^{p}_{u}(x_{0})$ involves a polynomial of degree at most $u$, so that $t^{p}_{u}(x_{0})$ is not a true subspace of $T^{p}_{\phi}(x_{0})$.
@@
whereas for every $t_{0}\in\mathbb{R}$, the Hölder exponent of $B$ at $t_{0}$ equals $h(t_{0})=\frac{1}{2}$. Therefore, if $\phi\in\mathcal{B}$ is such that $\phi(t)\sim(t\log\log 1/|t|)^{1/2}$ for all $t>0$, then $T^{\infty}_{\phi}$ is the adapted functional space that take into account the iterated logarithm in the pointwise regularity of $B$, a phenomenon that classical Hölder spaces fail to capture. In [12], they get pointwise estimates for solutions of elliptic partial differential equations in the context of $T^{p}_{\phi}$ spaces. They also get a more general version of Whitney’s extension Theorem.
@@ 3 Useful Results
It is the sufficient to prove the result for almost all $x\in A$. For $x\in A$, one has
@@
We are now in a position to generalize Rademacher’s Theorem. We make the proof of it [17] since we will need similar arguments for the new results.

### v3

@@ 2 Calderón-Zygmund spaces with Function parameter
For example, when $\underline{b}(\phi)$ is non-integer, the polynomial involved is of degree at most $\lfloor\underline{b}(\phi)\rfloor$. The condition $\underline{b}(\phi)>-\frac{d}{p}$ is here to ensure that the spaces are not degenerate. One can easily check that the polynomial appearing in (3) is unique. We set
@@
Then it is straightforward that $(T^{p}_{\phi}(x_{0}),\|\cdot\|_{T^{p}_{\phi}(x_{0})})$ is a Banach space. Let $t^{p}_{\phi,n}(x_{0})$, consisting of functions $f\in L^{p}(\mathbb{R}^{d})$ such that there exists a polynomial $P$ of degree at most $n$ such that
@@
If $\phi(t)=t^{u}$, with $u\in\mathbb{N}_{0}$, then $t^{p}_{\phi,u}(x_{0})$ is the usual space $t^{p}_{u}(x_{0})$ of [17]. Note that, in this case, the space $T^{p}_{\phi}(x_{0})$ involves a polynomial of degree at most u1u-1 whereas the space $t^{p}_{u}(x_{0})$ involves a polynomial of degree at most $u$, so that $t^{p}_{u}(x_{0})$ is not a true subspace of $T^{p}_{\phi}(x_{0})$.
@@
whereas for every $t_{0}\in\mathbb{R}$, the Hölder exponent of $B$ at $t_{0}$ equals $h(t_{0})=\frac{1}{2}$. Therefore, if $\phi\in\mathcal{B}$ is such that $\phi(t)\sim(t\log\log 1/|t|)^{1/2}$ for all $t>0$, then $T^{\infty}_{\phi}$ is the adapted functional space that take into account the iterated logarithm in the pointwise regularity of $B$, a phenomenon that classical Hölder spaces fail to capture.
The following result, proved in [12], is a function-parameter version of Whitney’s extension theorem. Roughly speaking, it shows that uniform $T^{p}_{\phi}$ conditions on a closed set $E$ ensure that the function coincides almost everywhere on $E$ with a $C^{n}$ function defined on a neighbourhood of $E$, together with quantitative control on finite differences of the derivatives of this $C^{n}$ extension. They have direct applications in the fine regularity theory with function parameters, in particular for obtaining pointwise regularity estimates for solutions of elliptic partial differential equations in the framework of $T^{p}_{\phi}$ spaces.
@@ 3 Useful Results
The following lemma is technically involved but plays a key role in the sequel. Its proof relies on a localization argument combined with covering techniques. Starting from the uniform control of the local $L^{p}$ averages of $f$ by $\phi(r)$, one reduces to a bounded setting where $f$ has compact support and identifies a large subset on which $f$ vanishes almost everywhere. Covering arguments then yield suitable integrability estimates for a singular kernel involving $\phi$, leading to the desired $o(\phi(r))$ decay for almost every $x\in E$ as $r\to 0^{+}$.
@@
It is sufficient to prove the result for almost all $x\in A$. For $x\in A$, one has
@@
We are now in a position to generalize Rademacher’s Theorem. We follow the proof in [17] since we will need similar arguments for the new results.